OpenStudy (anonymous):

3sqrt108 - 2sqrt18- 3sqrt48

OpenStudy (mertsj):

\[3\sqrt{108}-2\sqrt{18}-3\sqrt{48}=3\sqrt{36(3)}-2\sqrt{8(2)}-3\sqrt{16(3)}\]

OpenStudy (mertsj):

\[3(6)\sqrt{3}-2(3)\sqrt{2}-3(4)\sqrt{3}=18\sqrt{3}-6\sqrt{2}-12\sqrt{3}\]

OpenStudy (mertsj):

\[6\sqrt{3}-6\sqrt{2}\]

OpenStudy (mertsj):

I made a typo. The second radical should be 9(2) not 8(2) but I corrected it in the next step so the answer is still right.